Calculate the work done by a Force of 2𝑖 βˆ’ 𝑗 βˆ’ 3π‘˜ in moving an object from (2, 1, 3) to (9, 4, 6) where displacement is in metres.


1
Expert's answer
2021-06-08T11:56:04-0400
"\vec r=(9-2)\vec i+(4-1)\vec j+(6-3)\vec k"

"=7\vec i+3\vec j+3\vec k"


"W=\vec F\cdot \vec r=2(7)+(-1)(3)+(-3)(3)=2 (J)"

If the force is measured in newtons and the displacement is measured in metres, then the work is 2 J.